OpenStudy (bigdog):

solve (9/2x+6)-7/5x-15=2/3

OpenStudy (lukecrayonz):

x=290/93

OpenStudy (anonymous):

depends. if this is \[\frac{9}{2x+6}-\frac{7}{5x-15}=\frac{2}{3}\] then there is no answer

OpenStudy (bigdog):

how can you tell?